Almost anyone who visits Blackjacktournaments.com knows about Smith's quandary
of Doubling for Less when trailing as in Example(1) below shows.
Infinite Deck...................Last Hand..................Min Bet = 5.....................Max Bet = 500
Player..........BkRoll...................Bet...................Cards
BR2...............1000...............500 + 250..............2+3
BR1................1195.....................500.................10+10
Dealer....................................................................7
BR2 bets the max 500 and BR1 correlates by a matching bet of 500.
BR2 now doubles for 250.
The BR2 double is right in the middle of a double for 0 and a double for 500.
The BR2 double cannot be correlated by BR1 by splitting since he cannot split for less.
Actually any BR2 double of 200 up to 300 will put BR1 in the quandary.
If BR1 stands he gives up the high to BR2.
If BR1 splits he gives up the low to BR2. What should BR1 do?
Below are the probabilities of BR1 winning the session by the simple but reasonably
good strategy of splitting and playing not to bust against various dealer up cards.
Also given are the dealer bust probability with the various dealer up cards.
Dealer Up Card.............................................7..........................8......................10
Dealer Bust.............................................. .2623.................. .2447............... .2298
Prob of BR1 winning by
splitting and playing not to bust............ .8004................ .. .7818............... .7492
The best way for BR1 in Example(1) to play against the quandary with a dealer 7
up card is figured as follows:
The probability of BR1 winning by standing is BR1 needs the Dealer to make a hand.
Prob(Make a Hand) = 1 - Prob(Bust) = 1 - .2623 = .7377.
Since .7377 for standing is less than .8004 for splitting and playing not to bust;
then splitting and playing not to bust is the better way to go.
Instead of a dealer 7 up card in Example(1), what is BR1's best line of play against
an 8 or 10 dealer up card and each of their probabilities?
..............................................................BlueLight
as follows:
of Doubling for Less when trailing as in Example(1) below shows.
Infinite Deck...................Last Hand..................Min Bet = 5.....................Max Bet = 500
Player..........BkRoll...................Bet...................Cards
BR2...............1000...............500 + 250..............2+3
BR1................1195.....................500.................10+10
Dealer....................................................................7
BR2 bets the max 500 and BR1 correlates by a matching bet of 500.
BR2 now doubles for 250.
The BR2 double is right in the middle of a double for 0 and a double for 500.
The BR2 double cannot be correlated by BR1 by splitting since he cannot split for less.
Actually any BR2 double of 200 up to 300 will put BR1 in the quandary.
If BR1 stands he gives up the high to BR2.
If BR1 splits he gives up the low to BR2. What should BR1 do?
Below are the probabilities of BR1 winning the session by the simple but reasonably
good strategy of splitting and playing not to bust against various dealer up cards.
Also given are the dealer bust probability with the various dealer up cards.
Dealer Up Card.............................................7..........................8......................10
Dealer Bust.............................................. .2623.................. .2447............... .2298
Prob of BR1 winning by
splitting and playing not to bust............ .8004................ .. .7818............... .7492
The best way for BR1 in Example(1) to play against the quandary with a dealer 7
up card is figured as follows:
The probability of BR1 winning by standing is BR1 needs the Dealer to make a hand.
Prob(Make a Hand) = 1 - Prob(Bust) = 1 - .2623 = .7377.
Since .7377 for standing is less than .8004 for splitting and playing not to bust;
then splitting and playing not to bust is the better way to go.
Instead of a dealer 7 up card in Example(1), what is BR1's best line of play against
an 8 or 10 dealer up card and each of their probabilities?
..............................................................BlueLight
as follows: